# On the Mechanisms of Collaborative Learning in VAE Recommenders

### v1

@@ Abstract
Variational Autoencoders (VAEs) are a powerful alternative to matrix factorization for recommendation. A common technique in VAE-based collaborative filtering (CF) consists in applying binary input masking to user interaction vectors, which improves performance but remains underexplored theoretically. In this work, we analyze how collaboration arises in VAE‑based CF and show it is governed by latent proximity: we derive a latent sharing radius that informs when an SGD update on one user strictly reduces the loss on another user, with influence decaying as the latent Wasserstein distance increases. We further study the induced geometry: with clean inputs, VAE‑based CF primarily exploits local collaboration between input‑similar users and under‑utilizes global collaboration between far‑but‑related users. We compare two mechanisms that encourage global mixing and characterize their trade‑offs: ① $\beta$‑KL regularization directly tightens the information bottleneck, promoting posterior overlap but risking representational collapse if too large; ② input masking induces stochastic geometric contractions and expansions, which can bring distant users onto the same latent neighborhood but also introduce neighborhood drift. To preserve user identity while enabling global consistency, we propose an anchor regularizer that aligns user posteriors with item embeddings, stabilizing users under masking and facilitating signal sharing across related items. Our analyses are validated on the Netflix, MovieLens-20M, and Million Song datasets. We also successfully deployed our proposed algorithm on an Amazon streaming platform following a successful online experiment.
@@ 1 Introduction
Recommender systems are essential for delivering personalized user experiences. Having benefited from three decades of research (10.1145/192844.192905), Collaborative filtering (CF) remains an essential approach in today’s recommender systems (10.1145/3696410.3714908). CF predictive process consists in predicting user preferences by identifying and leveraging similarity patterns between users and items (su2009survey; ricci2010introduction), which naturally aligns with the framework of latent variable models (LVMs) (bishop2006pattern), where latent representations are used to capture the shared structure of user-item interactions. Due to their simplicity and effectiveness, LVMs have historically played a central role in CF research. However, these models are inherently linear, which limits their capacity to model the complex and non-linear nature of real-world user behavior (paterek2007improving; mnih2007probabilistic). To overcome these limitations, researchers have increasingly explored the integration of neural networks (NNs) into CF frameworks, enabling more expressive modeling and yielding notable improvements in recommendation accuracy (he2017neural; wu2016collaborative; liang2018variational; truong2021bilateral; li2021sinkhorn). A particularly successful line of work is VAE-based collaborative filtering (liang2018variational), which extends the variational autoencoder (VAE) framework (kingma2013auto; rezende2014stochastic) to collaborative filtering tasks. Unlike traditional latent factor models (hu2008collaborative; paterek2007improving; mnih2007probabilistic), which require learning a separate latent vector for each user, VAE-based models offer a user-independent parameterization, where the number of trainable parameters remains fixed regardless of the number of users lobel2019towards, leading to remarkable scalability. Additionally, empirically, VAE-based CF models consistently outperform many existing LVM-based alternatives (liang2018variational; kim2019enhancing; walker2022implicit; ma2019learning; guo2022topicvae; wang2023causal; wang2022disentangled; guo2024dualvae; husain2024geometric; li2021sinkhorn; tran2025varium).
@@
A key driver of the strong performance of VAE-based CF is the use of a binary mask that corrupts the user interaction vector, producing a partial history from which the model is trained to reconstruct the full interaction vector (liang2018variational). This masking strategy has been shown to significantly enhance recommendation accuracy (see Section 4.2 and Figure 2, comparing settings with and without masking). Although input noise has become a standard component in VAE-based CF models due to its empirical effectiveness, its underlying mechanisms and potential side effects remain largely unexplored. Existing works treat masking as a simple performance-enhancing heuristic, without thoroughly examining how it influences the learning process or affects the latent representations. Our work aims to fill this gap with a comprehensive study the effect of input noise in VAE-based CF. The contributions of this paper are summarized as follows:
@@ 2 Collaboration Mechanism in VAE-based CF
We index users by $u\in\{1,2,\ldots,U\}$ and items by $i\in\{1,2,\ldots,I\}$, where $U,I$ are the number of users and items, respectively. $\mathbf{X}\in\{0,1\}^{U\times I}$ represents the user-item interaction matrix (e.g., click, watch, check-in, etc.) and $\mathbf{x}_{u}=[\mathbf{x}_{u1},\mathbf{x}_{u2},\ldots,\mathbf{x}_{uI}]$ is an $I$-dimensional binary vector (the $u$-th row of $\mathbf{X}$) whereby $\mathbf{x}_{ui}=1$ implies that user $u$ has interacted with item $i$ and $\mathbf{x}_{ui}=0$ indicates otherwise. Note that $\mathbf{x}_{ui}=0$ does not necessarily mean user $n$ dislikes item $i$; the item may never be shown to the user. For simplicity, we use $\mathbf{x}$ to denote a general user interaction vector and retain $\mathbf{x}_{u}$ when specifically referring to user $u$. Additionally, we measure distances between input vectors with the $\ell_{1}$ norm $\|\cdot\|_{1}$ (Hamming) and between latent distributions with the $1$-Wasserstein distance $W_{1}(\cdot,\cdot)$. We refer to the Appendix Table A for the notation summary.
@@ 2.1 VAE-based Collaborative Filtering
Given a user’s interaction history $\mathbf{x}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{I}]^{\top}\in\{0,1\}^{I}$, the goal is to predict the full interaction behavior of this user with all remaining items. To simulate this process during training, as well as to avoid overfitting to non-informative patterns (steck2020autoencoders), VAE-based CF (liang2018variational; kim2019enhancing; truong2021bilateral; vanvcura2021deep) introduces a random binary mask $\mathbf{b}\in\{0,1\}^{I}$ is , with the entry 1 as un-masked, and 0 as masked. Thus, $\mathbf{x}_{h}=\mathbf{x}\odot\mathbf{b}$ is the user’s partial interaction history, and the goal is to recover the full $\mathbf{x}$ given $\mathbf{x}_{h}$.
@@ 2.2 Collaborative Learning via VAE-CF
Before presenting our theoretical analysis, we first recall the CF target: predict a user’s preferences from other users’ interaction patterns, without relying on item or user side content. Figure 1 illustrates two typical situations. First, users $\mathbf{x}_{1}$ and $\mathbf{x}_{4}$ are close in the input space (e.g., under $\ell_{1}$), so each serves as a natural reference for the other by the definition of the collaborative filtering. Second, let $S_{\mathbf{x}_{u}}=\{i:\ \mathbf{x}_{ui}=1\}$ contain the positive items for user $\mathbf{x}$, if $S_{\mathbf{x_{2}}}\subset S_{\mathbf{x_{1}}}$ i.e., $\mathbf{x}_{1}$ is a more active user with similar interests, then $\mathbf{x}_{1}$ should inform recommendations for $\mathbf{x}_{2}$ (e.g., items $i_{1},i_{2},i_{3}\in S_{\mathbf{x_{1}}}\setminus S_{\mathbf{x_{2}}}$). We refer to this latter, “far-but-related” influence as a global collaborative signal. To make these notions precise, for $\delta>0$, we define the input-space neighborhood for user $u$ $N_{\delta}(u)\;=\;\{\,v:\ \|\mathbf{x}_{u}-\mathbf{x}_{v}\|_{1}\leq\delta\,\}$ and the (nonzero) overlap indicator: $\mathrm{ov}(\mathbf{x}_{u},\mathbf{x}_{v})\;=\;\mathbf{1}_{\!\left\{\,\bigl|S_{\mathbf{x_{u}}}\cap S_{\mathbf{x_{v}}}\bigr|>0\,\right\}}.$
@@
A model exhibits a global collaborative signal at scale $\delta$ if there exist users $u$ and $v$ which are “far-but-related” i.e., $\|\mathbf{x}_{u}-\mathbf{x}_{v}\|_{1}>\delta$ and $\mathrm{ov}(\mathbf{x}_{u},\mathbf{x}_{v})=1$, and their predictions strictly influence each other.
We now examine how VAE-based CF methods leverage local and global collaborative signals.
@@
The quantity $D_{u,v}=L_{\theta z}\,W_{1}(q_{u},q_{v})+\Delta_{x}(u,v)$ is the transfer penalty: it upper-bounds the per-step regret at user $u$ by updating the model using user $v$. When $D_{u,v}<\|g_{u}\|$, the step computed on $v$ still impacts $u$; thus collaboration during training is localized to the latent neighborhood $\{\,v:\ W_{1}(q_{u},q_{v})<r_{\mathrm{share}}(u,v;\theta)\,\}$ and collaborative signal reduces with increased latent distance.
Consequently, to understand how VAE-CF exploits both local and global collaborative signals, we examine the correspondence between the geometry of the input space and the geometry of the latent space in VAE-based collaborative filtering, in the following section.
@@ 2.3 Impact of the Geometry of the Input and Latent Spaces
Additionally, assume the prior $p$ satisfies the Bobkov–Götze/Talagrand $T_{1}(C)$ inequality (bobkov1999exponential) with a constant $C>0$ (e.g., Normal prior $p=\mathcal{N}(0,\sigma^{2}I)$):
@@
  • Input locality induces latent locality. We show in Lemma 2.4 that nearby inputs (in $\ell_{1}$) induce nearby posteriors in $\mathcal{Z}$ i.e., input neighborhoods embed into latent neighborhoods, preserving local collaboration.
  • Reconstruction favors posterior separation. When $T(\mathbf{x}_{u})\neq T(\mathbf{x}_{v})$, any overlap of $q_{u}$ and $q_{v}$ incurs the strictly positive gap $\int(q_{u}+q_{v})\,\Delta_{A^{*}}\,d\mathbf{z}$ in Eq. (3), discouraging posterior overlap and preventing input‑distant, content‑mismatched users from becoming latent‑near, discouraging global collaboration.
  • $\beta$-KL encourages posterior overlap. Increasing $\beta$ reduces the optimal KL terms, tightening the upper bound of $W_{1}\!\big(q_{u},q_{v}\big)$, thereby increasing latent overlap i.e., potentially bringing input‑distant users closer in latent space and enabling global collaboration. However, an excessively large $\beta$ can lead to posterior collapse and weaken predictive performance. In practice, $\beta$ is typically kept small in ranking tasks.
Consequently, for moderate $\beta$, similar users remain latent‑near while content‑mismatched users are pushed apart, yielding a clustered latent geometry aligned with input similarity (Setting-1, Figure 2). By Theorem 2.3, SGD updates are therefore shared predominantly within these latent neighborhoods (i.e., across pairs with $W_{1}(q_{u},q_{v})<r_{\mathrm{share}}(u,v;\theta)$). Hence, with clean inputs, VAE‑CF primarily exploits local collaborative signals and suppresses global influence from input‑distant users.
@@
Random masks perturb pairwise distances stochastically i.e., each draw can contract or expand $\|\mathbf{x}^{\prime}_{u}-\mathbf{x}^{\prime}_{v}\|_{1}$. Under Lemma 2.4, contractions propagate to latent space; if the realized $\delta$ satisfies $L_{\phi}\,\delta<r_{\mathrm{share}}(u,v;\theta)$, an SGD step on $v$ strictly reduces $\mathcal{L}_{u}$ (cf. Theorem 2.3). On the other hand, expansions can push pairs outside the sharing radius, weakening local collaboration. This creates a pattern where latent representations of distant groups become intermittently entangled via contractions, while within-group cohesion diffuses due to occasional expansions (Setting-2, Figure 2).
@@ 3 Proposed Method: Personalized Item Alignment (PIA)
Prior work has primarily focused on the $\beta$-KL pathway i.e., adjusting $\beta$ or redesigning the prior $p(Z)$ (see Section B) while overlooking the issues introduced by the masking process. In this section, we explore an alternative direction for improving VAE-CF: addressing the downside of masking by stabilizing the stochastic geometry it induces, so that beneficial long-range sharing occurs more consistently and with reduced drift.
During training the encoder takes masked inputs $\mathbf{x}_{h}=\mathbf{x}\odot\mathbf{b}$ (cf. Eq. 2). Rather than conditioning $q_{\phi}(\mathbf{z}\mid\cdot)$ on auxiliary features at inference, we introduce a regularizer that, during training, pushes the posterior given $\mathbf{x}_{h}$ towards a user-specific target derived from the user’s positive items.
We define the overall objective as:
@@
Intuitively, this pulls the masked posterior toward the anchors of the items the user liked, making users with shared items close in the latent space.
@@
Proposition 3.1 indicates that PIA ① centers masked latents near the user’s item barycenter $\bar{\mathbf{e}}_{\mathbf{x}}$ and ② modestly reduces posterior spread. Two users $u,v$ with similar positive-item sets have close centroids, so their masked posteriors become latently near more frequently, increasing the chance they fall within the sharing radius and benefit from each other’s updates.

### v2

@@ Abstract
Variational Autoencoders (VAEs) are a powerful alternative to matrix factorization for recommendation. A common technique in VAE-based collaborative filtering (CF) consists in applying binary input masking to user interaction vectors, which improves performance but remains underexplored theoretically. In this work, we analyze how collaboration arises in VAE-based CF and show it is governed by latent proximity: we derive a latent sharing radius that informs when an SGD update on one user strictly reduces the loss on another user, with influence decaying as the latent Wasserstein distance increases. We further study the induced geometry: with clean inputs, VAE‑based CF primarily exploits local collaboration between input‑similar users and under‑utilizes global collaboration between far‑but‑related users. We compare two mechanisms that encourage global mixing and characterize their trade‑offs: ① $\beta$‑KL regularization directly tightens the information bottleneck, promoting posterior overlap but risking representational collapse if too large; ② input masking induces stochastic geometric contractions and expansions, which can bring distant users onto the same latent neighborhood but also introduce neighborhood drift. To preserve user identity while enabling global consistency, we propose an anchor regularizer that aligns user posteriors with item embeddings, stabilizing users under masking and facilitating signal sharing across related items. Our analyses are validated on the Netflix, MovieLens-20M, and Million Song datasets. We also successfully deployed our proposed algorithm on an Amazon streaming platform following a successful online experiment.
@@ 1 Introduction
Recommender systems are essential for delivering personalized user experiences. Having benefited from three decades of research (Resnick et al., 1994), Collaborative filtering (CF) remains an essential approach in today’s recommender systems (Zhu et al., 2025). CF predictive process consists in predicting user preferences by identifying and leveraging similarity patterns between users and items (Su and Khoshgoftaar, 2009; Ricci et al., 2010), which naturally aligns with the framework of latent variable models (LVMs) (Bishop and Nasrabadi, 2006), where latent representations are used to capture the shared structure of user-item interactions. Due to their simplicity and effectiveness, LVMs have historically played a central role in CF research. However, these models are inherently linear, which limits their capacity to model the complex and non-linear nature of real-world user behavior (Paterek, 2007; Mnih and Salakhutdinov, 2007). To overcome these limitations, researchers have increasingly explored the integration of neural networks (NNs) into CF frameworks, enabling more expressive modeling and yielding notable improvements in recommendation accuracy (He et al., 2017; Wu et al., 2016; Liang et al., 2018; Truong et al., 2021; Li et al., 2021). A particularly successful line of work is VAE-based collaborative filtering (Liang et al., 2018), which extends the variational autoencoder (VAE) framework (Kingma and Welling, 2013; Rezende et al., 2014) to collaborative filtering tasks. Unlike traditional latent factor models (Hu et al., 2008; Paterek, 2007; Mnih and Salakhutdinov, 2007), which require learning a separate latent vector for each user, VAE-based models offer a user-independent parameterization, where the number of trainable parameters remains fixed regardless of the number of users Lobel et al. (2019), leading to remarkable scalability. Additionally, empirically, VAE-based CF models consistently outperform many existing LVM-based alternatives (Liang et al., 2018; Kim and Suh, 2019; Walker et al., 2022; Ma et al., 2019; Guo et al., 2022; Wang et al., 2023; 2022; Guo et al., 2024; Husain and Monteil, 2024; Li et al., 2021; Tran and Lauw, 2025).
@@
A key driver of the strong performance of VAE-based CF is the use of a binary mask that corrupts the user interaction vector, producing a partial history from which the model is trained to reconstruct the full interaction vector (Liang et al., 2018). This masking strategy has been shown to significantly enhance recommendation accuracy (see Section 4.2 and Figure 2, comparing settings with and without masking). Although input noise has become a standard component in VAE-based CF models due to its empirical effectiveness, its underlying mechanisms and potential side effects remain largely unexplored. Existing works treat masking as a simple performance-enhancing heuristic, without thoroughly examining how it influences the learning process or affects the latent representations. Our work aims to fill this gap with a comprehensive study the effect of input noise in VAE-based CF. The contributions of this paper are summarized as follows:
@@ 2 Collaboration Mechanism in VAE-based CF
We index users by $u\in\{1,2,\ldots,U\}$ and items by $i\in\{1,2,\ldots,I\}$, where $U,I$ are the number of users and items, respectively. $\mathbf{X}\in\{0,1\}^{U\times I}$ represents the user-item interaction matrix (e.g., click, watch, check-in, etc.) and $\mathbf{x}_{u}=[\mathbf{x}_{u1},\mathbf{x}_{u2},\ldots,\mathbf{x}_{uI}]$ is an $I$-dimensional binary vector (the $u$-th row of $\mathbf{X}$) whereby $\mathbf{x}_{ui}=1$ implies that user $u$ has interacted with item $i$ and $\mathbf{x}_{ui}=0$ indicates otherwise. Note that $\mathbf{x}_{ui}=0$ does not necessarily mean user $u$ dislikes item $i$; the item may never be shown to the user. For simplicity, we use $\mathbf{x}$ to denote a general user interaction vector and retain $\mathbf{x}_{u}$ when specifically referring to user $u$. Additionally, we measure distances between input vectors with the $\ell_{1}$ norm $\|\cdot\|_{1}$ (Hamming) and between latent distributions with the $1$-Wasserstein distance $W_{1}(\cdot,\cdot)$. We refer to the Appendix Table A for the notation summary.
@@ 2.1 VAE-based Collaborative Filtering
Given a user’s interaction history $\mathbf{x}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{I}]^{\top}\in\{0,1\}^{I}$, the goal is to predict the full interaction behavior of this user with all remaining items. To simulate this process during training and to avoid overfitting to non-informative patterns (Steck, 2020), dropout-style random masking is commonly used in VAE-based CF (Wu et al., 2016; Liang et al., 2018; Lobel et al., 2019; Shenbin et al., 2020; Vančura and Kordík, 2021) as a form of stochastic input corruption. Concretely, we introduce a random binary mask $\mathbf{b}\in\{0,1\}^{I}$ is , with the entry 1 as un-masked, and 0 as masked. Thus, $\mathbf{x}_{h}=\mathbf{x}\odot\mathbf{b}$ is the user’s partial interaction history, the goal is to recover the full $\mathbf{x}$ given $\mathbf{x}_{h}$.
@@ 2.2 Collaborative Learning via VAE-CF
Before presenting our theoretical analysis, we briefly recall the goal of collaborative filtering (CF): to predict a user’s preferences from other users’ interaction patterns, without relying on user/item side information. Figure 1 highlights two common forms of cross-user information in CF.
(1) Neighborhood transfer. Users $\mathbf{x}_{1}$ and $\mathbf{x}_{4}$ are close in the input space (e.g., under $\ell_{1}$), so each serves as a natural reference for the other: similar interaction histories should lead to similar predictions.
(2) Far-but-related transfer. Users can be far in $\ell_{1}$ yet still meaningfully related through shared positives. Let $S_{\mathbf{x}_{u}}=\{i:\ \mathbf{x}_{ui}=1\}$ denote the set of positive items for user $u$. If $S_{\mathbf{x}_{2}}\subset S_{\mathbf{x}_{1}}$, then $\mathbf{x}_{1}$ is a more active user with similar interests and can inform recommendations for $\mathbf{x}_{2}$ (e.g., items $i_{1},i_{2},i_{3}\in S_{\mathbf{x}_{1}}\setminus S_{\mathbf{x}_{2}}$), even though $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ need not be close in $\ell_{1}$. We refer to this “far-but-related” influence as Global collaborative signal.
To formalize these notions, for $\delta>0$ we define the input-space neighborhood of user $u$ as
$N_{\delta}(u)\;=\;\{\,v:\ \|\mathbf{x}_{u}-\mathbf{x}_{v}\|_{1}\leq\delta\,\},$
and the (nonzero) overlap indicator as
$\mathrm{ov}(\mathbf{x}_{u},\mathbf{x}_{v})\;=\;\mathbf{1}_{\!\left\{\,\bigl|S_{\mathbf{x}_{u}}\cap S_{\mathbf{x}_{v}}\bigr|>0\,\right\}}.$
@@
A model exhibits a global collaborative signal at scale $\delta$ if there exist users $u$ and $v$ such that $\|\mathbf{x}_{u}-\mathbf{x}_{v}\|_{1}>\delta$ and $\mathrm{ov}(\mathbf{x}_{u},\mathbf{x}_{v})=1$, and their predictions strictly influence each other.
Cross-user influence via gradient transfer.
In VAE-CF, cross-user information is exchanged primarily during training because the same parameters are updated using many different users. We therefore operationalize “influence” as training-time transfer: user $v$ influences user $u$ if an SGD step computed from $v$ decreases $u$’s expected loss to first order. The following theorem formalizes this notion and shows that such transfer is governed by proximity of the users’ posteriors in latent space.
@@
Theorem 2.3 characterizes when an SGD step on user $v$ is beneficial for user $u$. The key quantity is the transfer penalty $D_{u,v}=L_{\theta z}\,W_{1}(q_{u},q_{v})+\Delta_{x}(u,v)$, which has two components: (i) a latent mismatch term $W_{1}(q_{u},q_{v})$ measuring how close the users’ posteriors are in latent space, and (ii) a content mismatch term $\Delta_{x}(u,v)$ capturing how different the decoder gradients are even when evaluated at the same latent code. To first order, a step on $v$ decreases $\mathcal{L}_{u}$ whenever $D_{u,v}<\|g_{u}(\theta)\|$. Therefore, gradient sharing is effectively restricted to a latent neighborhood around $u$, and the strength of collaboration decays as users become latently farther apart.
Consequently, Theorem 2.3 reduces the distinction between local and global collaborative signals (Definitions 2.12.2) to a geometric question: which user pairs become latently close enough to fall within the sharing radius. Local signals correspond to input-close pairs that remain latently close, whereas global signals require certain input-distant but overlap-related pairs to be mapped close in latent space. In the next section, we examine how the geometry of the input space (including masking) shapes the geometry of the latent space in VAE-based collaborative filtering.
@@ 2.3 Impact of the Geometry of the Input and Latent Spaces
Additionally, assume the prior $p$ satisfies the Bobkov–Götze/Talagrand $T_{1}(C)$ inequality (Bobkov and Götze, 1999) with a constant $C>0$ (e.g., Normal prior $p=\mathcal{N}(0,\sigma^{2}I)$):
@@
  • Local neighborhoods are preserved. Lemma 2.4 implies that $\ell_{1}$-near users induce nearby posteriors under $W_{1}$ (up to the encoder Lipschitz constant), so input-space locality naturally maps to latent-space locality, preserving local collaboration.
  • Reconstruction discourages overlap for mismatched users. Theorem 2.5 shows that when $T(\mathbf{x}_{u})\neq T(\mathbf{x}_{v})$, overlapping posteriors incur a strictly positive “compromise gap” $\int(q_{u}+q_{v})\,\Delta_{A^{*}}\,d\mathbf{z}$ in Eq. (3). This pushes the model to separate $q_{u}$ and $q_{v}$, making it difficult for input-distant, content-mismatched users to become latent-close, discouraging global collaboration.
  • $\beta$-KL encourages posterior overlap. Increasing $\beta$ reduces the optimal KL terms, tightening the upper bound of $W_{1}\!\big(q_{u},q_{v}\big)$, thereby increasing latent overlap i.e., potentially bringing input‑distant users closer in latent space and enabling global collaboration. However, overly strong KL pressure risks posterior collapse and weakens recommendation quality; in practice $\beta$ is typically kept small in ranking-focused CF.
Putting these pieces together, with clean (unmasked) inputs and moderate $\beta$, similar users remain latent‑close while content‑mismatched users are pushed apart, yielding a clustered latent geometry that mirror input similarity (Setting-1, Figure 2). By Theorem 2.3, SGD updates are therefore shared predominantly within these latent neighborhoods (i.e., across pairs with $W_{1}(q_{u},q_{v})<r_{\mathrm{share}}(u,v;\theta)$). In other words, under clean inputs, VAE‑CF primarily exploits local collaborative signals and suppresses global influence from input‑distant users.
@@
Theorem 2.6 quantifies how random masking can stochastically contract or expand the $\ell_{1}$ distance between two users’ interaction vectors. Under Lemma 2.4, contractions propagate to latent space; when the realized distance falls within the sharing radius (Theorem 2.3), an SGD step on $v$ benefits $u$, even if the original (unmasked) pair was far apart. This creates intermittent long-range sharing events that inject global collaborative signal. Conversely, expansions can push genuine neighbors outside the sharing radius, increasing gradient variance and weakening local transfer. The net effect is global mixing through occasional contractions at the cost of neighborhood drift from expansions (Setting-2, Figure 2).
@@ 3 Proposed Method: Personalized Item Alignment (PIA)
Prior work has primarily pursued the objective-level pathway (adjusting $\beta$ or redesigning $p(\mathbf{z})$; see Section B) while largely treating input masking as a benign training trick. In this section, we explore an alternative direction for improving VAE-CF: addressing the downside of masking by stabilizing the stochastic geometry it induces, so that beneficial long-range sharing occurs more consistently and with reduced drift. We introduce a training-only regularizer that pushes the masked posterior $q_{\phi}(\mathbf{z}\mid\mathbf{x}_{h})$ toward a user-specific target constructed from the user’s positive items. This stabilizes the masking-induced geometry without changing test-time inference, ensuring that different masked views of the same user map to a consistent latent region while bringing users with overlapping items closer in a semantically grounded way. We define the overall objective as:
@@
Proposition 3.1 indicates that PIA ① centers masked latents near the user’s item barycenter $\bar{\mathbf{e}}_{\mathbf{x}}$ and ② modestly reduces posterior spread. Two users $u,v$ with similar positive-item sets have close centroids, so their masked posteriors become latently close more frequently, increasing the chance they fall within the sharing radius and benefit from each other’s updates.
@@
Remark: geometric view of collaborative filtering.
The anchors $\mathbf{e}_{i}$} form a semantic map where co-liked items organize into neighborhoods, and $\bar{\mathbf{e}}_{\mathbf{x}}$ lies within the convex hull of the user’s liked items. This implies:
  • Users with large/diverse interaction histories. When $|S_{\mathbf{x}}|$ is large and diverse, $\bar{\mathbf{e}}_{\mathbf{x}}$ may be less peaked; PIA then primarily improves stability rather than enforcing a specific prototype. Crucially, users who share more items (or items from the same neighborhoods) have closer centroids, making their representations neighbors and strengthening collaborative signal via the sharing radius (Theorem 2.3). For multi-modal users, the barycenter lies between item neighborhoods, keeping them close to multiple communities; shared (or nearby) items still pull such users into overlapping regions, allowing global signals to flow from both sides. In effect, PIA brings similar users closer without collapsing diverse preferences into a single artificial mode.
  • Cold-start users. Warm users sculpt the semantic map: items they co-like are pulled into coherent neighborhoods. Cold users, even with few interactions, are placed into this already-organized map by aligning toward their clicked items’ anchors. Each positive item is a landmark surrounded by warm users who also liked it (and related items). Aligning cold users toward these landmarks means that even a single shared item can snap them into warm clusters. Moreover, because items sharing users are pulled closer under PIA, interacting with a similar item also places cold users in the right neighborhood. In click space, such users appear far from any warm user; in latent space, one or two well-placed items act as shortcuts to dense communities, enabling global collaborative signals to propagate to cold users.